OpenStudy (anonymous):

Last question for the night :D Consider the line that passes through the point (3, -6) and has a slope of 4. Part 1: Write the equation of this line using point-slope form. Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work.

ganeshie8 (ganeshie8):

you familiar with the 3 forms of line ? 1. point-slope form 2. slope-intercept form 3. standard form

OpenStudy (anonymous):

y - y1 = m(x-x1) Point slope form

OpenStudy (anonymous):

slope m = -6/3

OpenStudy (anonymous):

m = -2

OpenStudy (anonymous):

x1 = 3 y1 = -6 m = -2

OpenStudy (anonymous):

putting back inot original formula y - (-6) = -2(x-3) now rearranging y = -2(x-3) - 6 simplified y = -2x + 6 - 6 simplied further y = -2x

OpenStudy (anonymous):

http://openstudy.com/study# is that right?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

sorry how do u do the @ thing, get ganeshie to check it

OpenStudy (anonymous):

Okay write which part is which part like part 1: part 2: ...?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

how did u do that?

OpenStudy (anonymous):

just @ and write their name

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

hopefully he checks it. need it to be right. thanks for medal

ganeshie8 (ganeshie8):

looks good :)

ganeshie8 (ganeshie8):

oly one thing missing is, slope is given as 4 here . not -2. m = 4 :)

OpenStudy (anonymous):

yeah i know but if u do the problem it gives -2

ganeshie8 (ganeshie8):

hmm lets simply use the given slope :)

OpenStudy (anonymous):

m = rise / run right? m = -6/3

ganeshie8 (ganeshie8):

Part 1: Write the equation of this line using point-slope form. y - (-6) = 4(x-3)

ganeshie8 (ganeshie8):

Part 1 is done. Move to Part II.

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. y - (-6) = 4(x-3) y + 6 = 4x - 12 y = 4x - 18

ganeshie8 (ganeshie8):

Part 2 done. Move to Part III :)

OpenStudy (anonymous):

y = mx + b

ganeshie8 (ganeshie8):

^yes thats for Part II. I am sure you can do Part III. just rearrange the equation from Part II to standard form @InstagramDoll

OpenStudy (anonymous):

so u use y = 4x - 18 and put it into standard form is that right?

ganeshie8 (ganeshie8):

exactly.... :)

OpenStudy (anonymous):

which gets rearranged to: 4x - y - 18 = 0 is that right?

ganeshie8 (ganeshie8):

yes, take the 18 to other side :- Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. y = 4x - 18 4x - y = 18

ganeshie8 (ganeshie8):

Putting it all together, just for clarity... Make sure you understand this and work it on your own again @InstagramDoll Part 1: Write the equation of this line using point-slope form. y - y1 = m(x-x1) y - (-6) = 4(x-3) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. y - (-6) = 4(x-3) y + 6 = 4x - 12 y = 4x - 18 Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. y = 4x - 18 4x - y = 18

OpenStudy (anonymous):

ok thanks,

OpenStudy (anonymous):

I needa study all this tomorrow morning, lol i just wanted to quickly finish cause its 4:27 am and i needa be in bed. So the full answer would be . 1; y - (-6) = 4(x-3) 2; y - (-6) = 4(x-3) y + 6 = 4x - 12 y = 4x - 18 3; y = 4x - 18 4x - y = 18

OpenStudy (anonymous):

Thanks @ganeshie8 & @Helpmeplease99 for the help.

ganeshie8 (ganeshie8):

you certainlny deserve some good sleep.... after so much studying... good night :)

OpenStudy (anonymous):

Goodnight to you both too.